1/8+5/8=
I need help in calculating this problem step by step.

Answer by checkley79(3341)

1/8+5/8 with like denominators (8) you can just add the numerator values.
(1+5)/8=6/8 now reduce this fraction by dividing by 2.
6/8=3/4 ans.